# Certifying nonlocal properties of noisy quantum operations

### v1

@@ Abstract
We provide a unified framework to certify nonlocal properties of quantum channels from the correlations obtained in measurement protocols. Our approach gathers and extends fully- and semi-device independent methods for this purpose. We study the effect of different models of dephasing noise, some of which are shown to generate nonlocality and entanglement in special cases. In the extreme case of complete dephasing, the measurement protocols discussed yield particularly simple tests to certify nonlocality. These are based on the relations between bipartite quantum channels and their classical analogs: bipartite stochastic matrices defining conditional distributions.
I Introduction
Quantum theory predicts phenomena that cannot be explained with local hidden variable models [1, 2]. This property, known as nonlocality, is a resource for certain protocols in quantum information and computation [3, 4, 5]. These are implemented by using quantum states describing physical systems, which are processed and manipulated through quantum operations. Information is obtained by performing measurements, which provide a probability distribution in terms of relative frequencies of outcomes.
Therefore, it is of primal importance to identify nonlocal properties of both quantum states [6, 7] and the quantum operations acting on them [8, 9, 10], from the nonlocal behaviours in the outcome probability distributions [11, 12, 13]. Bell games are linear tests for measurement protocols that identify nonlocality in quantum states, by analyzing the correlations obtained through classical inputs and outputs [1, 2, 14]. By using quantum inputs, these can be modified to detect nonlocality of all entangled states [15]. Although these methods have been employed to study nonlocal properties of quantum states since several decades ago [1, 2, 15, 14], it was more recently that similar schemes were proposed to identify nonlocality in quantum channels [8, 9, 16, 10, 17, 18].
@@ I Introduction
In this paper we provide a framework to certify nonlocal properties of quantum channels, in the presence of dephasing noise. To this end, we use measurement protocols depicted in Fig 1 to formulate (semi-)device independent criteria to certify nonlocality of quantum channels. This approach includes and extends existing methods [8, 17] and allows us to identify the role of dephasing noise in the nonlocal properties of quantum channels: while local dephasing cannot generate nonlocality, there are cases in which a global dephasing operation can generate both entanglement and nonlocality. Our analysis in the case of complete decoherence yields simple criteria to detect nonlocality in quantum channels through a geometric insight into the transition from quantum to classical operations.
II Preliminaries
II.1 Quantum and classical states and channels
The set of quantum states $\rho$ acting on a Hilbert space of dimension $d$ is represented by $\mathcal{M}_{d}$. The set of classical states, namely the $d$-dimensional probability simplex, is denoted by $\Delta_{d}$. A classical channel transforming probability distributions is a stochastic matrix, that is, a matrix with non-negative entries that sum up to one in each column. The set of stochastic matrices with the dimension of inputs and outputs equal to $d$ is represented by $\mathcal{S}_{d}$. Analogously, completely positive (CP) and trace preserving (TP) maps, ${\cal E}_{A_{0}\rightarrow A_{1}}$, also known as quantum operations or quantum channels, transform quantum states. Their inputs and outputs are density matrices acting on a Hilbert space $\mathcal{H}_{A_{0}}$ and $\mathcal{H}_{A_{1}}$ with dimensions $d_{A_{0}}$ and $d_{A_{1}}$, respectively. When the input and output systems are held by the same party, we will denote by $A$ the joint system $A_{0}A_{1}$ and use the notation ${\cal E}_{A}$, or simply ${\cal E}$, if no confusion arises.
@@ II.1 Quantum and classical states and channels
The complete positivity and trace preserving (CPTP) conditions require that $J_{A_{0}A_{1}}\geq 0$ and
@@
Similarly, a classical superchannel $\Gamma$ [32], transforming stochastic matrices of size $d_{B_{0}}\times d_{A_{1}}$ to those of size $d_{B_{1}}\times d_{A_{0}}$, is isomorphic to a bipartite stochastic matrix of dimension $d_{A_{1}}d_{B_{1}}\times d_{A_{0}}d_{B_{0}}$. To be a valid superchannel, it has to additionally satisfy
@@
The set of quantum and classical superchannels acting on $d$-dimensional channels are respectively denoted by $\Theta_{d}$ and $\mathcal{T}_{d}$.
II.2 Nonlocality of bipartite distributions
@@
where $p_{\lambda}\geq 0$, $\sum_{\lambda}p_{\lambda}=1$. Otherwise, the distribution is called Bell nonlocal.
@@
for some shared state $\rho_{AB}$ and measurement effects $\{M^{a|x}\}$ and $\{N^{b|y}\}$. Distributions obtained from Eq. (7) are called quantum (Q).
@@
on the space of bipartite probability distributions, called Bell functionals. One denotes $\gamma_{S}:=\max_{p\in\mathcal{L}}\gamma(p)$ the maximum value of a Bell functional $\gamma$ over the set of distributions inside a selected subset $S$ (typically L, Q or NS). For example, the CHSH functional $\gamma_{ab,xy}=(-1)^{a+b}(-1)^{xy}$ satisfies $\gamma_{\operatorname{L{}}}=2$, $\gamma_{\operatorname{Q}}=2\sqrt{2}$ and $\gamma_{\operatorname{NS}}=4$.
II.3 Nonlocality in quantum states
A Bell protocol is a device independent local measurement protocol where a bipartite probability distribution is obtained from a bipartite quantum state through Eq. (7). Assuming quantum theory, the only subsets of bipartite distributions that can be obtained in this way are local (L) or quantum (Q). Separable quantum states can only produce local correlations, and therefore observing nonlocal quantum correlations is an indicator of the presence of entanglement [33, 34]. A quantum state $\rho_{AB}$ which cannot produce nonlocal correlations in the bipartition $A|B$ will be called a local state in such bipartition.
Remarkably, not all entangled states can produce correlations outside of L through Eq. (7), namely there exist local but entangled states [6, 35]. However, a subtle difference exists between separable and entangled states in certain measurement protocols: it was shown by Buscemi that all entangled states $\rho_{AB}$ produce correlations that cannot be obtained from separable states in a semi-device independent protocol with trusted quantum inputs [15],
@@ II.3 Nonlocality in quantum states
II.4 Nonlocality of bipartite channels
@@
where $T_{A}^{\lambda}\in\mathbb{R}^{d_{A_{1}}\times d_{A_{0}}}$ and $R_{B}^{\lambda}\in\mathbb{R}^{d_{B_{1}}\times d_{B_{0}}}$ are stochastic matrices. These stochastic matrices correspond with local probability distributions that decompose as Eq. (6), and thus we call them local stochastic matrices.
@@
II.5 The effect of dephasing noise in quantum states and channels
Here we will consider quantum noise producing decoherence through a dephasing channel $\mathcal{D}$, which leaves invariant the occupations of the input state in a preferred basis $\{\left|i\right\rangle\}$, $\left\langle i\right|\mathcal{D}(\rho)\left|i\right\rangle=\left\langle i\right|\rho\left|i\right\rangle$, while damping or inducing phases on the coherences. By imposing that a dephasing channel is a CPTP map, one verifies that it acts as [46, 47, 48, 49]
@@ II.5 The effect of dephasing noise in quantum states and channels
II.6 Complete decoherence: from quantum to classical
@@
where $\iota$ embeds a vector into the diagonal of a matrix. The CPTP condition on $\mathcal{E}_{A}$ implies that the vector $\operatorname{diag}(J^{\mathcal{E}})$ can be reshaped into a $d_{A_{1}}\times d_{A_{0}}$ stochastic matrix $S^{\mathcal{E}}$, which is a linear map between the $d_{A_{1}}$- and $d_{A_{0}}$-dimensional probability simplices. The stochastic map $S^{\mathcal{E}}$, known as classical action [54, 55, 49, 56, 57], will be called here the decoherent action of $\mathcal{E}$ [49] to emphasize that it corresponds to a possibly nonlocal channel arising from the effects of decoherence. The decoherent action can be obtained from Kraus operators $K_{i}$ of $\mathcal{E}$ through $S^{\mathcal{E}}=\sum_{i}K_{i}\odot K^{*}_{i}$, where * denotes the complex conjugation [58].
@@
The effect of complete decoherence on quantum sets discussed above is summarized in Fig. 5. These sets are obtained from the set $\mathcal{M}_{d^{4}}$ of four-partite states in different ways. Bipartite and monopartite states are obtained by partial trace. The probability simplices of corresponding sizes are obtained from these sets under decoherence. Bipartite and monopartite quantum channels are obtained from four-partite and bipartite states through the cross-section of Eq. (2). In particular, a generic quantum state $\sigma_{A_{0}B_{0}A_{1}B_{1}}$ with full-rank gives a generic quantum channel through $J_{A_{0}B_{0}A_{1}B_{1}}=Y\sigma_{A_{0}B_{0}A_{1}B_{1}}Y$ where $X=\mathrm{Tr}_{A_{0}A_{1}}\sigma_{A_{0}B_{0}A_{1}B_{1}}$ and $Y=\mathbf{1}\otimes 1/\sqrt{X}$ imposes the partial trace condition [59]. Then a classical map (stochastic matrix) is obtained due to decoherence as in Eq. (19) from the diagonal entries of the Choi matrix $J$ describing the bipartite channel. The transition from channels to superchannels can be seen as an extended Choi-Jamiołkowski isomorphism, as superchannels are isomorphic to one-way quantum nonsignaling channels through Eq. (4c). These decohere to one-way nonsignalling distributions through Eq. (5).
@@
Thus one identifies classical subsets embedded in quantum subsets at different levels: general distributions, conditional distributions and nonsignaling distributions are embedded in quantum states, channels and superchannels. In this way, recent studies on the geometry of the sets of local and nonlocal correlations [60, 22] are directly linked to the research on geometry of quantum states [58, 61] and geometry of quantum entanglement [62, 63]. Furthermore, analysis of the relative volume of the set of bipartite local correlations [64, 65, 66] can be compared with the problem of estimating the relative volume of the set of bipartite separable states [67, 68, 69, 70, 71].
III Channel nonlocality from local measurements
@@
Figure 1: Schemes to obtain bipartite conditional probability distributions (bipartite stochastic matrices) from bipartite channels: (a) A shared entangled state is processed with local operations (with ancillary systems) to produce the inputs which undergo the channel ${\cal E}$ to be analyzed, and then local measurements are performed to obtained the probabilities of Eq. (20). (b) Product states prepared locally undergo a bipartite channel and are then local gneralized measurements are performed to obtain the distribution in Eq. (26). This case can be seen as bipartite distributions being obtained after local superchannels are performed in the form of $\Xi_{A}\otimes\Xi_{B}[{\cal E}_{AB}]$. (c) A bipartite distribution in Eq. (39) is obtained with inputs in product states spanning a preferred basis, which undergo a bipartite channel and are then projective-measured in the same preferred basis.
III.1 Local measurement protocols with shared entanglement
@@
Note that Proposition 1 provides a criterion to detect quantum channels outside of the sets LOSR, LOSE and qNS which is fully device independent, in the sense that none of the quantum devices (the shared state $\rho$, the input channels $\Lambda^{x}$ and $\Phi^{y}$, and the measurements $M^{a}$ and $N^{b}$) is trusted.
Figure 2: Subsets of distributions obtained from bipartite operations. Left. Inclusions of the following subsets of quantum channels (CPTP): quantum nonsignaling (qNS), local operations assisted by the shared resources of entanglement (LOSE) and randomness (LOSR). Right. Inclusions of the subsets of the polytope of bipartite stochastic matrices ($\mathcal{S}$): nonsignaling (NS), quantum (Q) and local (L) correlations. Bell functionals $\gamma$ bounding each subset of bipartite distributions bound the corresponding subsets of bipartite channels, through Prop. 1, Cor. 2 and Prop. 7. Here we depict the linear cuts delimited by CHSH functional.
III.2 Local measurement protocols without entanglement

### v3

@@ Abstract
Certifying quantum properties from the probability distributions they induce is an important task for several purposes. While this framework has been largely explored and used for quantum states, its extrapolation to the level of channels started recently in a variety of approaches. In particular, little is known about to what extent noise can spoil certification methods for channels.
In this work we provide a unified methodology to certify nonlocal properties of quantum channels from the correlations obtained in prepare-and-measurement protocols: our approach gathers fully and semi-device-independent existing methods for this purpose, and extends them to new certification criteria. In addition, the effect of different models of dephasing noise is analysed. Some noise models are shown to generate nonlocality and entanglement in special cases. In the extreme case of complete dephasing, the measurement protocols discussed yield particularly simple tests to certify nonlocality, which can be obtained from known criteria by fixing the dephasing basis. These are based on the relations between bipartite quantum channels and their classical analogues: bipartite stochastic matrices defining conditional distributions.
1 Introduction
Quantum theory predicts phenomena that cannot be explained with local hidden-variable models [1, 2]. This property, known as nonlocality, is a resource for certain protocols in quantum information and computation [3, 4]. These are implemented by using quantum states describing physical systems, which are processed and manipulated through quantum operations. Information is obtained by performing measurements, which provide a probability distribution in terms of relative frequencies of outcomes.
Therefore, it is of primal importance to identify nonlocal properties of both quantum states [5, 6] and the quantum operations acting on them [7, 8, 9], from the nonlocal behaviors in the outcome probability distributions [10, 11, 12]. Bell games are linear tests for measurement protocols that identify nonlocality in quantum states, by analysing the correlations obtained through classical inputs and outputs [1, 2, 13]. By using quantum inputs, these can be modified to detect nonlocality of all entangled states [14]. Although these methods have been employed to study nonlocal properties of quantum states for several decades [1, 2, 14, 13], it was more recently that similar schemes were proposed to identify nonlocality in quantum channels [7, 8, 15, 9, 16, 17].
@@ 1 Introduction
In this paper we provide a framework to certify nonlocal properties of quantum channels, in the presence of dephasing noise. To this end, we use measurement protocols depicted in Fig. 1 to formulate (semi-)device-independent criteria to certify nonlocality of quantum channels. This approach includes and extends existing methods [7, 16] and allows us to identify the role of dephasing noise in the nonlocal properties of quantum channels: while local dephasing cannot generate nonlocality, there are cases in which a global dephasing operation can generate both entanglement and nonlocality. Our analysis in the case of complete decoherence yields simple criteria to detect nonlocality in quantum channels through a geometric insight into the transition from quantum to classical operations.
2 Preliminaries
2.1 Quantum and classical states and channels
The set of quantum states $\rho$ acting on a Hilbert space of dimension $d$ is represented by $\mathcal{M}_{d}$. The set of classical states, namely the $d$-dimensional probability simplex, is denoted by $\Delta_{d}$. A classical channel transforming probability distributions is a stochastic matrix, that is, a matrix with non-negative entries that sum up to one in each column. The set of stochastic matrices with the dimension of inputs and outputs equal to $d$ is represented by $\mathcal{S}_{d}$. Analogously, completely positive (CP) and trace-preserving (TP) linear maps, ${\cal E}_{A_{0}\rightarrow A_{1}}$, also known as quantum operations or quantum channels [27, §10.3], transform quantum states. Their inputs and outputs are density matrices acting on a Hilbert space $\mathcal{H}_{A_{0}}$ and $\mathcal{H}_{A_{1}}$ with dimensions $d_{A_{0}}$ and $d_{A_{1}}$, respectively. When the input and output systems are held by the same party, we will denote by $A$ the joint system $A_{0}A_{1}$ and use the notation ${\cal E}_{A}$, or simply ${\cal E}$, if no confusion arises.
@@ 2.1 Quantum and classical states and channels
Here, $\mathbf{1}_{X}$ denotes the identity operator on $\mathcal{H}_{X}$ and the superscript $T$ the transpose operation. The complete positivity and trace preserving (CPTP) conditions require that $J_{A_{0}A_{1}}\geq 0$ and
@@
Similarly, a classical superchannel $\Gamma$ [32], transforming stochastic matrices of size $d_{B_{0}}\times d_{A_{1}}$ to those of size $d_{B_{1}}\times d_{A_{0}}$, corresponds with a bipartite stochastic matrix of dimension $d_{A_{1}}d_{B_{1}}\times d_{A_{0}}d_{B_{0}}$. To be a valid superchannel, it has to additionally satisfy
@@
Physically, equations  (4c) and (5) impose nonsignaling conditions: the outputs cannot signal to the inputs; see also Eqs. (15a) and (15b). The set of quantum and classical superchannels acting on $d$-dimensional channels are respectively denoted by $\Theta_{d}$ and $\mathcal{T}_{d}$.
2.2 Nonlocality of bipartite distributions
@@
where $p_{\lambda}\geq 0$ and $\sum_{\lambda}p_{\lambda}=1$. Otherwise, the distribution is called Bell nonlocal.
@@
for some shared state $\rho_{AB}$ and positive operator-valued measurements (POVM) $M^{x}$ and $N^{y}$. That is, $M^{a|x},N^{b|y}\geq 0$ for each effect and $\sum_{a}M^{a|x}=\mathbf{1}_{A}$, $\sum_{b}N^{b|y}=\mathbf{1}_{B}$ [27, §10.1]. Distributions obtained from Eq. (7) are called quantum (Q).
@@
on the space of bipartite probability distributions, called Bell functionals. One denotes $\gamma_{S}:=\max_{p\in S}\gamma(p)$ the maximum value of a Bell functional $\gamma$ over the set of distributions inside a selected subset $S$ (typically L, Q or NS). For example, the Clauser-Horne-Shimony-Holt (CHSH) [33] functional $\gamma_{ab,xy}=(-1)^{a+b}(-1)^{xy}$ satisfies $\gamma_{\operatorname{L{}}}=2$, $\gamma_{\operatorname{Q}}=2\sqrt{2}$ and $\gamma_{\operatorname{NS}}=4$ [34, 13].
2.3 Nonlocality in quantum states
A Bell protocol is a device-independent local measurement protocol where a bipartite probability distribution is obtained from a bipartite quantum state through Eq. (7). Assuming quantum theory, the only subsets of bipartite distributions that can be obtained in this way are local (L) or quantum (Q). Separable quantum states can only produce local correlations, and therefore observing nonlocal quantum correlations is an indicator of the presence of entanglement [35, 36]. A quantum state $\rho_{AB}$ which cannot produce nonlocal correlations in the bipartition $A|B$ will be called a local state in such bipartition.
Remarkably, not all entangled states can produce correlations outside of L through Eq. (7), namely there exist local but entangled states [5, 37]. However, a subtle difference exists between separable and entangled states in certain measurement protocols: it was shown by Buscemi that all entangled states $\rho_{AB}$ produce correlations that cannot be obtained from separable states in a semi-device-independent protocol with trusted quantum inputs [14],
@@ 2.3 Nonlocality in quantum states
2.4 Nonlocality of bipartite channels
@@
where $T_{A}^{\lambda}\in\mathds{R}^{d_{A_{1}}\times d_{A_{0}}}$ and $R_{B}^{\lambda}\in\mathds{R}^{d_{B_{1}}\times d_{B_{0}}}$ are stochastic matrices. These stochastic matrices correspond with local probability distributions that decompose as Eq. (6), and thus we call them local stochastic matrices.
@@
2.5 The effect of dephasing noise in quantum states and channels
Here we will consider quantum noise producing decoherence through a dephasing channel $\mathcal{D}$, which leaves invariant the occupations of the input state in a preferred basis $\{\left|i\right\rangle\}$, $\left\langle i\right|\mathcal{D}(\rho)\left|i\right\rangle=\left\langle i\right|\rho\left|i\right\rangle$, while damping or inducing phases on the coherences. By imposing that a dephasing channel is a CPTP linear map, one verifies that it acts as [48, 49, 50, 51]
@@ 2.5 The effect of dephasing noise in quantum states and channels
2.6 Complete decoherence: from quantum to classical
@@
where $\iota$ embeds a vector into the diagonal of a matrix. The CPTP conditions on $\mathcal{E}_{A}$ imply that the vector $\operatorname{diag}(J^{\mathcal{E}})$ can be reshaped into a $d_{A_{1}}\times d_{A_{0}}$ stochastic matrix $S^{\mathcal{E}}$, which is a linear map between the $d_{A_{1}}$- and $d_{A_{0}}$-dimensional probability simplices. The stochastic map $S^{\mathcal{E}}$, known as classical action [56, 57, 51, 58, 59], will be called here the decoherent action of $\mathcal{E}$ [51] to emphasize that it corresponds to a possibly nonlocal channel arising from the effects of decoherence. The decoherent action can be obtained from Kraus operators $K_{i}$ of $\mathcal{E}$ through $S^{\mathcal{E}}=\sum_{i}K_{i}\odot K^{*}_{i}$, where * denotes the complex conjugation [60].
@@
The effect of complete decoherence on quantum sets discussed above is summarized in Fig. 5. These sets are obtained from the set $\mathcal{M}_{d^{4}}$ of four-partite states in different ways. Bipartite and monopartite states are obtained by partial trace. The probability simplices of corresponding sizes are obtained from these sets under decoherence. Bipartite and monopartite quantum channels are obtained from four-partite and bipartite states through the cross-section of Eq. (2). In particular, a generic quantum state $\sigma_{A_{0}B_{0}A_{1}B_{1}}$ with full-rank gives a generic quantum channel through $J_{A_{0}B_{0}A_{1}B_{1}}=Y\sigma_{A_{0}B_{0}A_{1}B_{1}}Y$ where $X=\mathrm{Tr}_{A_{0}A_{1}}\sigma_{A_{0}B_{0}A_{1}B_{1}}$ and $Y=\mathbf{1}\otimes 1/\sqrt{X}$ imposes the partial trace condition [61]. Then a classical map (stochastic matrix) is obtained due to decoherence as in Eq. (19) from the diagonal entries of the Choi matrix $J$ describing the bipartite channel. The transition from channels to superchannels can be seen as an extended Choi-Jamiołkowski isomorphism, as superchannels are isomorphic to one-way quantum nonsignaling channels through Eq. (4c). These decohere to one-way nonsignaling distributions through Eq. (5).
@@
Thus one identifies classical subsets embedded in quantum subsets at different levels: general distributions, conditional distributions and nonsignaling distributions are embedded in quantum states, channels and superchannels. In this way, recent studies on the geometry of the sets of local and nonlocal correlations [62, 21] are directly linked to the research on geometry of quantum states [60, 27, 63] and geometry of quantum entanglement [64, 65]. Furthermore, analysis of the relative volume of the set of bipartite local correlations [66, 67, 68] can be compared with the problem of estimating the relative volume of the set of bipartite separable states [69, 70, 71, 72, 73].
3 Channel nonlocality from local measurements
@@
Figure 1: Schemes to obtain bipartite conditional probability distributions (bipartite stochastic matrices) from bipartite channels: (a) A shared entangled state is processed with local operations (with ancillary systems) to produce the inputs which undergo the channel ${\cal E}$ to be analysed, and then local measurements are performed to obtained the probabilities of Eq. (20). (b) Product states prepared locally undergo a bipartite channel and then local measurements are performed to obtain the distribution in Eq. (26). This case can be seen as bipartite distributions being obtained after local superchannels are performed in the form of $\Xi_{A}\otimes\Xi_{B}[{\cal E}_{AB}]$. (c) A bipartite distribution in Eq. (39) is obtained with inputs in product states spanning a preferred basis, which undergo a bipartite channel and are then projective-measured in the same preferred basis.
3.1 Local measurement protocols with shared entanglement
@@
Here, $\text{id}_{RS}$ denotes the identity channel in $RS$, $\mathcal{E}$, $\Lambda^{x}$ and $\Phi^{y}$ generic quantum channels, $M^{a}$ and $N^{b}$ POVMs and $\rho$ a quantum state. The subindices indicate their respective systems. The following generalizes existing results in detecting non-LOSR quantum operations [7, 16], allowing for nonproduct input states –Fig. 1 (a).
@@
Note that Proposition 1 provides a criterion to detect quantum channels outside of the sets LOSR, LOSE and QNS which is fully device-independent, in the sense that none of the quantum devices (the shared state $\rho$, the input channels $\Lambda^{x}$ and $\Phi^{y}$, and the measurements $M^{a}$ and $N^{b}$) is trusted.
Figure 2: Subsets of distributions obtained from bipartite operations. Left. Inclusions of the following subsets of quantum channels (CPTP): quantum nonsignaling (QNS), local operations assisted by the shared resources of entanglement (LOSE) and randomness (LOSR). Right. Inclusions of the subsets of the polytope of bipartite stochastic matrices ($\mathcal{S}$): nonsignaling (NS), quantum (Q) and local (L) correlations. Bell functionals $\gamma$ bounding each subset of bipartite distributions bound the corresponding subsets of bipartite channels, through Prop. 1, Cor. 2 and Prop. 7. Here we depict the linear cuts delimited by some Bell functional.
3.2 Local measurement protocols without entanglement